# Operator Growth in Disordered Spin Chains: Indications for the Absence of Many-Body Localization

### v3

@@ Abstract
We consider the growth of a local operator $A$ in one-dimensional many-body systems with Hamiltonian $H$ by calculating the $k$-fold commutator $[H,[H,[...,[H,A]]]]$. We derive general bounds for the operator norm of this commutator in free and interacting fermionic systems with and without disorder thus directly connecting the operator growth hypothesis with questions of localization. We show, in particular, that in a localized system the norm does grow at most exponentially and that the contributions of operators to the total norm are exponentially suppressed with their length. We support our general results by considering one specific example, the XXZ chain with random magnetic fields. We solve the operator spreading in the XX case without disorder exactly. For the Anderson and Aubry-André models we provide strict upper bounds. We support our results by symbolic calculations of the commutator up to high orders. For the XXX case with random magnetic fields, these symbolic calculations show a growth of the operator norm faster than exponential and consistent with the general bound for an ergodic system. Also, there is no exponential decay of the contribution of operators as function of their length. We conclude that there is no indication for a many-body localization transition. Finally, we also discuss the differences between the interacting and non-interacting cases when trying to perturbatively transform the microscopic to an effective Hamiltonian of local conserved charges by consecutive Schrieffer-Wolff transformations. We find that such an approach is not well-defined in the interacting case because the transformation generates $\sim 4^{\ell}$ terms connecting sites a distance $\ell$ apart which can overwhelm the exponential decay with $\ell$ of the amplitude of each individual term.
@@ I Introduction
Studying the possibility of localization in the many-body case has gained renewed interest in the last 17 years starting with the work by Basko, Aleiner, and Altshuler [5]. This was followed by a large number of numerical studies concentrating, in particular, on the Heisenberg (XXX) chain with random magnetic fields [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. By a Jordan-Wigner transform, this model is equivalent to a fermionic chain with nearest-neighbor hoppings, density-density interactions, and random on-site potentials. Based on an analysis of spectral properties such as the level spacing as well as of dynamical properties such as the growth of the entanglement entropy after a quantum quench from an unentangled initial state, an apparent consensus was reached that this model shows a phase transition in all eigenstates at a finite disorder strength, separating an ergodic from a many-body localized phase. This consensus was first called into question by studies showing that the transition point, defined for a finite system of size $L$, appears to slowly shift to infinite disorder strength for $L\to\infty$ [20, 21]. These arguments were supported by studies of the number entropy after a quantum quench [22, 23, 24, 25, 26, 27] and of the fidelity susceptibility [28, 29].
Here we want to propose and follow a new avenue by studying the properties of operators instead of wave functions. Our motivation is the following observation: In an Anderson localized model, the projectors onto the eigenstates $P_{n}=|E_{n}\rangle\langle E_{n}|$ are conserved quasi-local operators, i.e., they are centered at a site in the lattice and decay exponentially away from this site. These operators are also often called local integrals of motion (LIOMs). Without loss of generality, we can order the operators such that $P_{n}$ is an operator localized around lattice site $n$. To be more precise, we can use the fermionic language in which the Anderson Hamiltonian and any unitary transformation of this Hamiltonian are bilinear. We can then write $P_{n}=\sum_{ij}a^{n}_{ij}c^{\dagger}_{i}c_{j}$ where $c_{i}^{(\dagger)}$ is a fermionic annihilation (creation) operator at site $i$ and $|a^{n}_{ij}|\sim\exp(|i-n|)\exp(|j-n|)$ are amplitudes. In particular, the contribution of bilinear operators to $P_{n}$ is exponentially suppressed with the distance they cover.
There is a unitary transformation $U$ which diagonalizes the microscopic, non-interacting Anderson Hamiltonian $H$, bringing it into the form $\tilde{H}=UHU^{-1}=\sum_{n}E_{n}P_{n}$ where $P_{n}$ is quasi-local and conserved, $[H,P_{n}]=0$. Importantly, this unitary transformation itself is quasi-local and will always map local microscopic operators $A_{i}$ onto quasi-local operators, $\tilde{A}_{i}=UA_{i}U^{-1}$. As we will show, this has important implications for the spreading of a local operator in Euclidean time. We can write
@@
I.e., the spreading of the operator is encoded in the $k$-fold commutator $[H,A_{i}]^{(k)}$ with the Hamiltonian. It is important to note that the Frobenius norm of this commutator is not affected by the unitary transformation $||[H,A_{i}]^{(k)}||_{2}=||[\tilde{H},\tilde{A}_{i}]^{(k)}||_{2}$. For the latter we expect that the norm of operators in the commutator which cover a distance of $l$ sites is exponentially suppressed with $l$ at fixed order $k$. As we will show, this implies that the total norm of all terms in the commutator can only grow exponentially with $k$.
@@
Here, $J_{ij}$ are exponentially decaying coupling constants which describe a slow dephasing between the conserved charges. For the Anderson case, $J_{ij}=0$ and the $\tau^{z}_{n}$ can be chosen to be the projection operators $P_{n}$. We note that the choice of the quasi-local conserved operators $\tau^{z}_{n}$ is not unique. We will argue in the following that in the interacting case a local operator in the microscopic model should also remain local, up to exponential tails, under Euclidean time evolution as in the non-interacting case. This implies, in particular, that the total norm in both cases should at most grow exponentially with $k$.
This has to be contrasted with the general, ergodic case. In one dimension, there is no phase transition in a clean system at finite temperatures. Thus, the time evolution is always analytic and Eq. (I) implies that the norm of the $k$-fold commutator $[H,A_{i}]^{(k)}$ has to grow slower than factorially. In Ref. [30], it has been shown that this is indeed the case, however, the derived bound for the total norm grows much faster than exponential and almost factorially, $||[H,A_{i}]^{(k)}||_{2}\sim(k/\ln k)^{k}$. The operator growth hypothesis states that any generic quantum chaotic system will asymptotically saturate this bound [30, 31]. Thus, we have a sharp contrast between the growth of the operator norm of the $k$-fold commutator of a local operator with the Hamiltonian between a localized, non-ergodic system, where it can grow at most exponentially, and a non-localized system, where an almost factorial growth is expected in general. This qualitative difference in the operator norm growth is what we will investigate here. To avoid misunderstandings, we want to stress already here that what we mean by a local operator and what we will investigate in the following is an operator $A_{i}$ which resides on lattice site $i$ in the microscopic model. In contrast, operators of the form $A=\sum_{i}A_{i}$ are also often called local conserved charges if $[H,A]=0$ and are of particular importance in integrable models. Operators of the type $A$, which are sums of local densities but which extend over the entire lattice, are also the type of operators whose growth has been investigated in Refs. [31, 32]. The latter type of operators are not of interest for a study of localization and we will not consider them here.
@@
Here we want to expand the study of operator growth to disordered systems. There is, in particular, a direct conflict between the operator growth hypothesis which suggests that an almost factorial growth is indicative of ergodicity and systems such as the random transverse Ising model which has been suggested to have a many-body localized phase but does always show almost factorial operator growth [37]. While we will consider Lieb-Robinson bounds and the decay of off-diagonal matrix elements in App. B and App. C, it is important to note that these quantities are obtained by taking sums over $k$-fold commutators $[H,A_{i}]^{(k)}$ and thus some of the information, in particular about the spatial support of such commutators which is relevant for the question of localization, is lost. Here we follow a novel path by directly obtaining analytical results for the asymptotic scaling of $k$-fold commutators resolved by the extent of their spatial support. Our results resolve the above mentioned conflict and suggest that the operator growth hypothesis is valid and that many-body systems, even for strong disorder, do not localize but rather stay ergodic.

### v4

@@ Abstract
We consider the spreading of a local operator $A$ in one-dimensional many-body systems with Hamiltonian $H$ by calculating the $k$-fold commutator $[H,[H,[...,[H,A]]]]$. We derive bounds for the operator norm of this commutator in free and interacting systems with and without disorder thus directly connecting the operator growth hypothesis with questions of localization. We analytically show, in particular, that an almost factorial growth of the operator norm—as recently proven for the random Ising model and strongly suggested to also hold for the Heisenberg model with random fields here—is inconsistent with an exponential localization of $A$. Assuming that a quasi-local unitary $U$ exists which maps $H$ onto an effective Hamiltonian $\tilde{H}=UHU^{\dagger}=\sum_{n}E_{n}\tau^{z}_{n}+\sum_{i,j}J_{ij}\tau^{z}_{i}\tau^{z}_{j}+\dots$, we show that $\tilde{A}=UAU^{\dagger}$ is a quasi-local operator which in the many-body case, in contrast to the Anderson case, indeed does not remain exponentially localized in general leading to an almost factorial growth of the commutator norm. Therefore the unitary $U$ in many-body systems with maximal norm growth either does not exist and such systems are always ergodic or unusual non-ergodic phases described by $\tilde{H}$ do exist which violate the operator growth hypothesis and in which local operators spread over the entire lattice implying that transport will eventually set in. To investigate this issue further, we concentrate on the XXZ chain with random magnetic fields. We analytically and symbolically verify our general results for the non-interacting Anderson and Aubry-André models. For the XXX case, the symbolic calculations are consistent with a maximal norm growth. Furthermore, we find no indication of a weakened exponential localization of $A$, expected for strong disorder and low commutator orders if the unitary $U$ does exist. Finally, we study the differences between the interacting and non-interacting cases when trying to perturbatively construct $U$ by consecutive Schrieffer-Wolff transformations. While it is straightforward to show that this construction converges in the Anderson case, we find no indications for a convergence in the interacting case, suggesting that $U$ does not exist and that many-body localization is absent.
@@ I Introduction
Studying the possibility of localization in the many-body case has gained renewed interest in the last 17 years starting with the work by Basko, Aleiner, and Altshuler [5]. This was followed by a large number of numerical studies concentrating, in particular, on the Heisenberg (XXX) chain with random magnetic fields [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19]. By a Jordan-Wigner transform, this model is equivalent to a fermionic chain with nearest-neighbor hoppings, density-density interactions, and random on-site potentials. Based on an analysis of spectral properties such as the level spacing as well as of dynamical properties such as the growth of the entanglement entropy after a quantum quench from an unentangled initial state, an apparent consensus was reached that this model shows a phase transition in all eigenstates at a finite disorder strength, separating an ergodic from a many-body localized phase. This consensus was first called into question by studies showing that the transition point, defined for a finite system of size $L$, appears to slowly shift to infinite disorder strength for $L\to\infty$ [20, 21]. These arguments were supported by studies of the number and Hartley entropies after a quantum quench [22, 23, 24, 25, 26, 27] and of the fidelity susceptibility [28, 29].
Here we want to propose and follow a new avenue by studying the properties of operators instead of wave functions which offer a unified perspective of the non-interacting and interacting cases. Our motivation is the following observation: In an Anderson localized model, the projectors onto the eigenstates $P_{n}=|E_{n}\rangle\langle E_{n}|$ are conserved quasi-local operators, i.e., they are centered at a site in the lattice and decay exponentially away from this site. These operators are also often called local integrals of motion (LIOMs). Without loss of generality, we can order the operators such that $P_{n}$ is an operator localized around lattice site $n$. To be more precise, we can use the fermionic language in which the Anderson Hamiltonian and any unitary transformation of this Hamiltonian are bilinear. We can then write $P_{n}=\sum_{ij}a^{n}_{ij}c^{\dagger}_{i}c_{j}$ where $c_{i}^{(\dagger)}$ is a fermionic annihilation (creation) operator at site $i$ and $|a^{n}_{ij}|\sim\exp(|i-n|)\exp(|j-n|)$ are amplitudes. In particular, the contribution of bilinear operators to $P_{n}$ is exponentially suppressed with the distance they cover.
There is a unitary transformation $U$ which diagonalizes the microscopic, non-interacting Anderson Hamiltonian $H$, bringing it into the form $\tilde{H}=UHU^{\dagger}=\sum_{n}E_{n}P_{n}$ where $P_{n}$ is quasi-local and conserved, $[H,P_{n}]=0$. Importantly, this unitary transformation itself is quasi-local and will always map local microscopic operators $A_{i}$ onto quasi-local operators, $\tilde{A}_{i}=UA_{i}U^{\dagger}$. As we will show, this has important implications for the spreading of a local operator in Euclidean time. We can write
@@
I.e., the spreading of the operator is encoded in the $k$-fold commutator $[H,A_{i}]^{(k)}$ with the Hamiltonian. It is important to note that the Frobenius norm of this commutator is not affected by the unitary transformation $||[H,A_{i}]^{(k)}||_{2}=||[\tilde{H},\tilde{A}_{i}]^{(k)}||_{2}$. We also note that the same $k$-fold commutator needs to be considered when calculating real-time two-point correlation functions as, for example, a current-current correlation function which determines the transport properties of a system.
@@
Here, $J_{ij}$ are exponentially decaying coupling constants which describe a slow dephasing between the conserved charges. For the Anderson case, $J_{ij}=0$ and the $\tau^{z}_{n}$ can be chosen to be the projection operators $P_{n}$. We note that the choice of the quasi-local conserved operators $\tau^{z}_{n}$ is not unique. If we define a many-body localized phase by the property that a local operator in the microscopic model remains local—up to exponential tails—under Euclidean time evolution as in the non-interacting case, then this implies, as we will show, that the total operator norm in both cases can at most grow exponentially with the commutator order $k$. We will also show that if local operators spread over the entire system then this strongly suggests that transport will eventually set in.
This has to be contrasted with the general, ergodic case. In one dimension, there is no phase transition in a clean system at finite temperatures. Thus, the time evolution is always analytic and Eq. (I) implies that the norm of the $k$-fold commutator $[H,A_{i}]^{(k)}$ has to grow slower than factorially. In Ref. [30], it has been shown that this is indeed the case; however, the derived bound for the total norm grows much faster than exponentially and almost factorially: $||[H,A_{i}]^{(k)}||_{2}\sim(k/\ln k)^{k}$. The operator growth hypothesis states that any generic quantum chaotic system will asymptotically saturate this bound [30, 31]. Thus, we have a sharp contrast between the growth of the operator norm of the $k$-fold commutator of a local operator with the Hamiltonian between a strictly exponentially localized, non-ergodic system, where it can grow at most exponentially, and a non-localized system, where an almost-factorial growth is expected in general. This qualitative difference in the operator norm growth is what we will investigate here. To avoid misunderstandings, we want to stress already here that what we mean by a local operator and what we will investigate in the following is an operator $A_{i}$ which resides on lattice site $i$ in the microscopic model. In contrast, operators of the form $A=\sum_{i}A_{i}$ are also often called local conserved charges if $[H,A]=0$ and are of particular importance in integrable models. Operators of the type $A$, which are sums of local densities but which extend over the entire lattice, are also the type of operators whose growth has been investigated in Refs. [31, 32]. The latter type of operators are not of interest for a study of localization and we will not consider them here. From hereon we will therefore drop the lattice index with the understanding that $A$ is always a local operator.
@@
Here we want to expand the study of operator growth to disordered systems. There is, in particular, a direct conflict between the operator growth hypothesis which suggests that an almost factorial growth is indicative of ergodicity and systems such as the random transverse Ising model, which has been suggested to have a many-body localized phase but always shows almost factorial operator growth independent of the disorder strength [37]. While we will consider Lieb-Robinson bounds and the decay of off-diagonal matrix elements in the Suppl. Mat. [38], it is important to note that these quantities are obtained by taking sums over $k$-fold commutators $[H,A]^{(k)}$ and thus some of the information, in particular about the spatial support of such commutators which is relevant for the question of localization, is lost. Here we follow a novel path by directly obtaining analytical results for the asymptotic scaling of $k$-fold commutators resolved by the extent of their spatial support. Our results are a step towards resolving the above mentioned conflict and suggest that the operator growth hypothesis is valid and that many-body systems, even for strong disorder, do not localize but rather stay ergodic.
@@
II Outline and main results
Here we want to give a summary of our main results and provide a guide to allow the reader to find these results and to point out parts of the paper which can be skipped in a first reading.
It has recently come to light that in the Ising model with random transverse and longitudinal fields [49, 48]
$||[H,\sigma^{z}_{0}]^{(k)}||_{2}\sim\left(\frac{k}{\ln k}\right)^{k}$ (II.1)
is both an upper and a lower bound [37]. This asymptotically maximal growth suggests, according to the operator growth hypothesis, that this model is ergodic for all disorder strengths. In Sec. V we show data from symbolic calculations which indicate that the same maximal growth also holds for the Heisenberg model with random fields. This means that either both models do not show an MBL phase or that the operator growth hypothesis has to be modified such that a maximal operator growth is also possible in certain non-ergodic models.
To investigate this issue further, we resolve the commutator by the number of sites $l$ it covers. At commutator order $k$, the commutator $[H,A]^{(k)}$ of a local operator $A$ acting on a single site with a nearest-neighbor Hamiltonian contains operators which span at most $l=k+1$ sites. We show in Sec. III.1 that the contribution of operators with length $l$ to the total norm is always bounded by
$||[H,A]_{l}^{(k)}||_{2}\leq(2CJ)^{k}2^{l}S(k,l)$ (II.2)
where $CJ$ is a constant and $S(k,l)$ the Stirling number of the second kind. This bound, which is expected to be asymptotically tight for an ergodic system, has a maximum at some $l(k)$ which shifts to infinity for $k\to\infty$. I.e., the operator spreads through the entire lattice. Summing over all operator lengths $l$ leads to the almost factorial bound (II.1). On the other hand, we might expect that in a strictly exponentially localized system, the operator $A$ remains quasi-local, $||[H,A]_{l}^{(k)}||_{2}\sim\text{e}^{-l}$. In this case the total norm growth is dominated by terms with support on a few sites only. This can be understood in more detail by representing the commutator as graphs which we present in Sec. V.1 for the non-interacting case and in Sec. V.2 for the interacting case. In the strictly exponentially localized case we then find that
$||[H,A]_{l}^{(k)}||_{2}\sim\text{e}^{k}\text{e}^{-l}\,.$ (II.3)
This implies, in particular, that in the strictly exponentially localized case the total norm cannot grow faster than exponentially with the commutator order $k$. Conversely, a faster, almost factorial growth of the total norm as in the many-body random-field Ising and Heisenberg models is incompatible with an exponential localization of operators. This important result is summarized in Fig. 1.
Figure 1: If local operators remain strictly exponentially localized then the total norm cannot grow faster than exponential. Conversely, in many-body systems such as the Ising and Heisenberg models with random fields the almost factorial growth of the total norm implies that local operators will not remain exponentially localized.
In the non-interacting case, we prove that in a localized phase, local operators indeed remain strictly exponentially localized. In Sec. V.1 we demonstrate, in particular, that in an Anderson localized phase Eq. (II.3) holds while in a non-localized phase $||[H,A]_{l}^{(k)}||_{2}\sim\text{e}^{k}$ independent of $l$ for $k\gg l$. I.e., in the latter case the operator spreads through the entire lattice. We show, furthermore, that this difference in scaling allows one to unambiguously identify the phase transition in the Aubry-André model. We do find it important to rigorously establish this novel operator perspective on localization by studying the non-interacting case in detail first. However, the reader who does accept the non-interacting results and is mostly interested in what this approach has to say about MBL might skip Sec. V.1 in its entirety.
If one takes strict exponential localization of local operators under commutation with the Hamiltonian as the very definition of a localized phase, then one has to conclude that the Ising model with random transverse and longitudinal fields and also—based on the analysis presented in this paper—the Heisenberg model with random fields do not possess a localized phase because in these two models the total norm growth is faster than exponential and described by Eq. (II.1), independent of the disorder strength.
On the other hand, if one takes the existence of a unitary transformation $U$ which maps these microscopic models onto an effective model (I.2) as the definition then there appears to be potentially room for a faster than exponential growth of the total norm in the interacting case as shown by the strict upper bounds derived in Sec. IV. Here we summarize the main results. First, it is important to note that the Frobenius norm is invariant under unitary transformations $||[H,A]^{(k)}||_{2}=||[\tilde{H},\tilde{A}]^{(k)}||_{2}$ with $\tilde{H}=UHU^{\dagger}$ and $\tilde{A}=UAU^{\dagger}$. It is easy to show that for a local operator such as $\sigma^{x}_{0}$ we have $||[\tilde{H},\sigma^{x}_{0}]^{(k)}||_{2}\sim\text{e}^{k}\text{e}^{-l}$ with $\tilde{H}$ as given in Eq. (I.2). I.e., the operator growth hypothesis does hold for $\tilde{H}$ clearly indicating that the model is not ergodic. However, it is crucial to note that a local operator $A$ in the microscopic model gets mapped onto a quasi-local operator $\tilde{A}$ under the unitary transformation $U$. In the non-interacting case, this is still a one-body operator and we can show that this operator remains strictly exponentially localized. I.e., in the non-interacting case, the definitions of a localized phase by requiring that $A$ stays strictly exponentially localized and the alternative definition that a quasi-local unitary $U$ exists which maps $H$ to $\tilde{H}$ are equivalent.
In the interacting case, however, $\tilde{A}$ is in general a generic quasi-local operator and, if no additional microscopic restrictions exist, then $||[\tilde{H},\tilde{A}]^{(k)}||_{2}$ will show an almost factorial growth. Furthermore, we will demonstrate in Sec. IV that we can obtain the following upper bound for the spreading of a quasi-local operator
$||[\tilde{H},\tilde{A}]_{l}^{(k)}||_{2}\lesssim l^{k}\text{e}^{-\kappa l}$ (II.4)
where $\kappa=\kappa(D)$ increases with disorder strength $D$. We argue in Sec. IV that $||[H,A]_{l}^{(k)}||_{2}$ should show the same functional dependence as Eq. (II.4) for strong disorder. We note that this function has a maximum at $l=k/\kappa$, i.e., for large commutator order $k$ the local operator $A$ spreads over the entire lattice. Furthermore, for very large $k$ the bound (II.2) will eventually be tighter, implying that the operator will spread through the lattice in the exact same way as in an ergodic system. This then leads to two possible scenarios: (1) The quasi-local unitary $U$ does exist and the many-body Hamiltonian $H$ has an unusual non-ergodic phase where it is equivalent to $\tilde{H}$ but shows maximal norm growth and local operators $A$ delocalize. We show in Sec. IV.3 that this strongly suggests that transport eventually sets in at long times. I.e., even if such a phase exists it should not be called a many-body localized phase. (2) The unitary $U$ does not exist. These results are summarized in Fig. 2.
Figure 2: Assuming that a quasi-local unitary $U$ with $\tilde{H}=UHU^{\dagger}$ exists, local operators $A$ are mapped onto quasi-local operators $\tilde{A}=UAU^{\dagger}$. In the free-fermion case, $\tilde{A}$ remains a one-body operator, implying that $A$ is strictly exponentially localized and the total norm growth is exponential. In contrast, $\tilde{A}$ is a generic quasi-local operator in the many-body case in general, implying that $A$ eventually delocalizes and that the total norm growth is almost factorial.
One way to investigate this question further is by symbolic calculations of $||[H,A]_{l}^{(k)}||_{2}$ which, according to Eq. II.4, should be dominated by an exponential decay for strong disorder and finite commutator order $k$. Our results presented in Sec. V.2 are not inconsistent with Eq. (II.4), however, even for disorder strengths much larger than the hopping amplitude, $D\gg J$, the exponential decay is never fully dominant. We therefore also study by symbolic calculations in Sec. VI whether a mapping $H\to\tilde{H}$ by successive Schrieffer-Wolff transformations is convergent. This has been suggested to be possible by Imbrie [48] for the Ising model with random fields and is supposed to work similarly to the Anderson case. However, there are important differences: (i) In the Anderson case, one-body operators remain one-body operators under Schrieffer-Wolff transformations. This means that two sites which are a distance $\ell$ apart do only get coupled by $c_{j}^{\dagger}c_{j+\ell}+h.c.$ and that this coupling is suppressed by $(J/D)^{2\ell}$ if we ignore resonances. In the interacting case, on the other hand, there is no such restriction and we can place any of the operators $\{\mathbbm{I},\sigma^{x},\sigma^{y},\sigma^{z}\}$ on the sites $j+1,\cdots,j+\ell-1$ and a Pauli operator on sites $j,\ell$ leading to $3\cdot 4^{\ell-1}$ different operators coupling site $j$ and site $\ell$. Even if each individual term is suppressed as $D^{-2\ell}$ the parameter for the total contribution $\sim 4^{\ell}D^{-2\ell}$ is no longer always small. This is, of course, not necessarily inconsistent with having a transition at some finite $D$. (ii) The smallest pairwise difference of random fields on a chain segment of length $\ell$ scales as $\ell^{-2}$. This implies that the effective coupling scales as $J_{j,j+\ell}\sim D^{-2\ell}\ell^{2}$ in the non-interacting case and resonances therefore do not proliferate. However, in the interacting case we have many-body resonances caused by some linear combination of all the random fields on the segment. We show in Fig. 18 results of symbolic Schrieffer-Wolff transformations which do not show any evidence that the amplitudes of the generated long-range terms do converge. (iii) Finally, we note that for a finite system we want to stop the Schrieffer-Wolff transformations once the norm of the off-diagonal parts of the effective Hamiltonian falls below some threshold $\varepsilon$. For the non-interacting case this leads to the condition that $k\sim\ln(L/\varepsilon)$ steps are required for a chain of length $L$ which is fully consistent with our symbolic calculations. In the interacting case, on the other hand, the number of generated off-diagonal terms scales exponentially with system size. If we assume that each individual contribution is still suppressed as $D^{-2k}$, i.e. resonances do not proliferate, then this leads to the condition $k\sim(L-\ln\varepsilon)/2\ln D$. In particular, we now expect a linear scaling with system size, $k\sim L$, as compared to the logarithmic scaling in the non-interacting case. However, we find symbolically that the Schrieffer-Wolff transformations for typical disorder distributions do not converge in $k\sim L$ steps. Instead, our results appear to be consistent with $k\sim\text{e}^{L}$, i.e., all the terms generated will need to be eliminated. This suggests that many-body resonances prohibit the perturbative construction of an effective Hamiltonian via Schrieffer-Wolff transformations in the thermodynamic limit.